Prove that

(i) the three points A(1, 4, 2), B(3, 2, 4) and C(5, 0, 6) are collinear,

(ii) the four points P(2, ;1, 1), Q(1, 3, ;2), R((2, 1, ;3) and S(3, 2, 0) are coplanar.



1
Expert's answer
2021-02-02T01:02:37-0500

(i)

"A(1, 4, 2), B(3, 2, 4) and C(5, 0, 6)"

The three point A,B,C are collinear if the direction ratio of AB,BC,AC are proportional.

Now,

"A(1, 4, 2), B(3, 2, 4)"

Direction Ratio "=3-1,2-4,4-2"

"= \n(2,-2,2)"

"so, a_1=2,b_1=-2,c_1=2"

Now,

"B(3, 2, 4) , C(5, 0, 6)"

Direction Ratio"=" "(\n5-3,0-2,6-4)"

"=(2,-2,2)"

So,

"a_2=2,b_2=-2,c_2=2"

ratios,

"a_2\/a_1=2\/2=1"

"b_2\/b_1=-2\/-2=1"

"c_2\/c_1=2\/2=1"


Therefore A,B,C are collinear

Hence proved.


(ii)

"P(2, 1, 1), Q(1, 3, 2), R(2, 1, 3) and S(3, 2, 0)"

So, let

"\overrightarrow{OP}=2{\hat{i} +\hat{j}}+\hat{k}"

"\overrightarrow{OQ}={\hat{i} +3\hat{j}}+2\hat{k}"

"\overrightarrow{OR}=2{\hat{i} +\hat{j}}+3\hat{k}"

"\overrightarrow{OS}={\hat{i} +2\hat{j}}"

now,

"\overrightarrow{PQ}=OQ-OP={\hat{i} +2\hat{j}}+\hat{k}"

"\overrightarrow{PR}=OE-OP=2\hat{k}"

"\overrightarrow{PS}={\hat{i} +\hat{j}}-\hat{k}"

Solving aboove three equations we get

"|PQ,PR,PS|=0"

Hence the given four points are coplanar.

Hence proved .